# seq-JEPA: Autoregressive Predictive Learning of Invariant-Equivariant World Models

### v2

@@ Abstract
Current self-supervised algorithms commonly rely on transformations such as data augmentation and masking to learn visual representations. This is achieved by enforcing invariance or equivariance with respect to these transformations after encoding two views of an image. This dominant two-view paradigm often limits the flexibility of learned representations for downstream adaptation by creating performance trade-offs between high-level invariance-demanding tasks such as image classification and more fine-grained equivariance-related tasks. In this work, we proposes seq-JEPA, a world modeling framework that introduces architectural inductive biases into joint-embedding predictive architectures to resolve this trade-off. Without relying on dual equivariance predictors or loss terms, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to specified transformations and another invariant to them. To do so, our model processes short sequences of different views (observations) of inputs. Each encoded view is concatenated with an embedding of the relative transformation (action) that produces the next observation in the sequence. These view-action pairs are passed through a transformer encoder that outputs an aggregate representation. A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation. Empirically, seq-JEPA demonstrates strong performance on both equivariant and invariant benchmarks without sacrificing one for the other. Furthermore, it excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
@@ 1 Introduction
Equivariance is a crucial representational property for downstream tasks that require fine-grained distinctions. For example, given representations that are invariant to color, it is not possible to distinguish between certain species of flowers or birds (Lee et al., 2021; Xiao et al., 2021). Moreover, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024). While some equivariant SSL approaches have reported minor gains on tasks typically associated with invariance (e.g., classification) (Devillers and Lefort, 2022; Park et al., 2022; Gupta et al., 2023), a growing body of work highlights a fundamental trade-off between learning invariance and equivariance, i.e., models that capture equivariance-related style latents do not fare well in classification and vice versa (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024; Rusak et al., 2025). This trade-off has recently received theoretical support (Wang et al., 2024), underscoring the need for new architectural and objective designs that can reconcile these competing goals.
@@
Our results demonstrate that individual encoded views in seq-JEPA become transformation/action-equivariant. Through ablations, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network. In contrast, the aggregate representation of views, produced at the output of the transformer, becomes largely transformation/action-invariant. This emergent architectural disentanglement of invariant and equivariant representations is central to seq-JEPA’s competitive performance compared to invariant and equivariant SSL methods on both categories of tasks (Figure 3). Unlike most prior equivariant SSL methods (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), our model does not rely on explicitly crafted loss terms or objectives to achieve equivariance, nor is it instructed to learn the decomposition of two representation types. Instead, the dual representation structure arises naturally from the model architecture and action-conditioned predictive learning.
Beyond resolving the invariance-equivariance trade-off, seq-JEPA further benefits from moving away from the two-view paradigm in SSL. We show that it performs well on tasks requiring integration over sequences of observations. In one scenario, inspired by embodied vision in primates, our model learns image representations without augmentations or masking, solely by predicting across simulated eye movements (saccades). In another setting, it performs path integration over sequences of actions—such as eye movements or 3D object rotations in 3DIEBench (Garrido et al., 2023). Our key contributions are as follows:
  • We introduce seq-JEPA, a self-supervised world model that learns architecturally distinct equivariant and invariant representations through sequential prediction over action-observation pairs, without requiring explicit equivariance losses or dual predictors.
@@ 2.2 Architecture
A backbone encoder, $f$ encodes the first $M$ views, producing representations $\{z_{i}\}_{i=1}^{M}$. Except for $z_{M}$, each $z_{i}$ is concatenated with its corresponding action embedding and passed to a transformer encoder $g$ (no MLP projector is used after the encoder), which aggregates the sequence of action-observation pairs. The transformer uses a learnable [AGG] token (analogous to the [CLS] token in ViTs (Dosovitskiy et al., 2020)) to generate the aggregate representation:
@@
The ground truth $z_{M+1}$ is computed using a target encoder—an exponential moving average (EMA) of $f$. The target representation is passed through a stop-gradient operator (sg) to avoid representational collapse. The training objective is to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$:
@@
Figure 1: seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn architecturally distinct invariant and equivariant representations.
@@ 2.3 Action and observation sets
To evaluate generalization across transformation types, we define three sets of action-observation pairs (Figure 2). See Appendix A.1 for details of each setup.
3D Invariant Equivariant Benchmark (3DIEBench). We first evaluate on the 3DIEBench dataset (Garrido et al., 2023), designed for equivariance analysis. It includes 3D object renderings with variations in rotation, floor hue, and lighting. We primarily study equivariance to $SO(3)$ rotations and secondarily to appearance factors of floor and light hue.
Hand-Crafted Augmentations. In this setting, we use transformed views generated via common SSL augmentations (e.g., crop, color jitter, blur), where actions correspond to relative augmentation parameters. We use CIFAR100 and Tiny ImageNet, and follow EquiMod’s augmentation protocol (Devillers and Lefort, 2022).
Predictive Learning Across Saccades (PLS). Going beyond conventional transformations such as augmentations or 3D rotations, we show that seq-JEPA can learn visual representations from a sequence of partial observations thanks to architectural inductive biases—without relying on any hand-crafted augmentations. Our PLS has a similar flavor to I-JEPA (Assran et al., 2023) but does not require engineered masking strategies. In PLS, we train seq-JEPA on sequences of patches extracted from high-resolution images. For instance, with STL-10, we use $32\times 32$ patches to form the observation sequence. In this setting, actions correspond to the relative positions between patch centers, simulating saccadic eye movements and inducing 2-D positional equivariance to representations. To select fixation points, we adopt two biologically inspired techniques that increase informativeness, reduce redundancy, and improve the downstream utility of the aggregate representation (Figure 2):
@@ 3.1 Compared methods and baselines
We compare seq-JEPA against both invariant and equivariant SSL baselines. Invariant methods include SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022). Equivariant methods include SEN (Park et al., 2022), EquiMod (Devillers and Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024). For all baselines, architectural details are given in Appendix A.3. We also evaluate two hybrid baselines based on our architecture:
@@ 3.2 Training protocol
All models use ResNet-18 (He et al., 2016) as the backbone encoder. For action conditioning, we use a learnable linear projection to learn action embeddings (default action embedding is 128-d). The sequence aggregator in seq-JEPA is a lightweight transformer encoder (Vaswani et al., 2017) with three layers and four attention heads. The predictor is a 2-layer MLP with 1024 hidden units and ReLU activation. In order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train baselines that typically use MLP projectors in two variants: (1) with original MLP projector; and (2) with the MLP replaced by a transformer encoder and a sequence length of one. We report the better-performing variant to ensure a fair comparison. All models are trained from scratch with a batch size of 512. We use 1000 epochs for 3DIEBench and 2000 epochs for other datasets to obtain asymptotic performance. We use AdamW for models with transformer projectors (including seq-JEPA) due to its stability and improved regularization in transformer training. For ConvNet-only models with MLP heads, we use the Adam optimizer. Full hyperparameters are detailed in Appendix A.
@@ 3.3 Evaluation metrics and protocol
To assess equivariance, we follow the protocol of Garrido et al. (2023) and train a regressor on frozen encoder representations to predict the relative transformation (action) between two views. We report the $R^{2}$ score on the test set. As a proxy measure of invariance, we use top-1 classification accuracy of a linear probe on top of frozen representations. For all baselines, probes are trained on encoder outputs. For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 1) and report the number of observation views used during training and inference. For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 1) for seq-JEPA models in Appendix B.4. Training details of evaluation heads are given in Appendix A.2.
@@ 4.1 Quantitative evaluation on 3DIEBench
We use the 3DIEBench benchmark to quantitatively compare invariant and equivariant representations in seq-JEPA with baseline methods. This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object classification. Table 3 provides a summary of our evaluation on the 3DIEBench where equivariant methods have been conditioned on rotation. In addition to the relative rotation between two views, in the last column we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view. For seq-JEPA, we trained models with varying training sequence lengths (denoted by $M_{tr}$ in the table) and measure the linear classification performance on top of aggregate representations with different inference lengths (denoted by $M_{val}$).
Among invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance. Among equivariant baselines, SIE and ContextSSL yield strong rotation prediction performance due to their specialized equivariance predictors and loss functions, but underperform in classification. EquiMod and SEN offer better classification performance, yet compromise equivariance. In contrast, seq-JEPA achieves strong performance in both, matching the best rotation $R^{2}$ scores while exceeding baselines in classification, with gains increasing with inference sequence length $M_{val}$. Ablating action conditioning leads to a sharp drop in equivariance but retains classification accuracy, confirming our hypothesis that action-conditioned sequential aggregation enables a segregated invariant-equivariant representation structure. For additional results on 3DIEBench including models with varying training and inference sequence lengths and models conditioned on both rotation and color, see Appendices B.1 to B.3.
@@
Table 1: Evaluation on 3DIEBench for linear probe classification (invariance) and rotation prediction (equivariance). Equivariant models and seq-JEPA are conditioned on rotation. For seq-JEPA, training and inference sequence lengths are denoted by $M_{tr}$ and $M_{val}$.
@@ 4.3 Evaluation with Hand-Crafted Augmentations
We assess invariance and equivariance under hand-crafted augmentations by training on CIFAR100 and Tiny ImageNet (Table 2). Models with action conditioning are trained conditioned on crop, color jitter, blur, or all three (indicated in the first column of the table). seq-JEPA consistently achieves higher equivariance than both invariant and equivariant baselines across all transformations. Notably, except for the model trained on CIFAR-100 and conditioned on blur, the best equivariance performance for a given augmentation is achieved when the model is specialized and conditioned only on that augmentation. Furthermore, ablating actions (last row in the table) causes seq-JEPA to lose its equivariance across transformations compared to action-conditioned models. Overall, our model outperforms both invariant and equivariant families in terms of equivariance, while being competitive in terms of classification performance. For additional results with varying training and inference sequence lengths, see Appendix B.3.
@@ 4.4 Predictive Learning across Saccades and Path Integration
In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations to learn visual representations in a predictive manner. Table 6 shows that seq-JEPA reaches 83.44% top-1 accuracy on STL-10, comparable to SimCLR (85.23%) trained with full-resolution images and strong augmentations. This gap narrows further when increasing the inference sequence length from $M_{val}=4$ to $M_{val}=6$. Ablating action conditioning causes the accuracy on top of the aggregate representations to drop sharply, indicating that 2-D positional awareness is essential to forming semantic representations across simulated eye movements. Compared to Conv-JEPA, which accumulates prediction losses pairwise, seq-JEPA performs better in classification, highlighting the importance of sequence aggregation when dealing with partial observations in SSL. Further ablations show that saliency-driven sampling and IoR are critical for forming informative, non-overlapping patch sequences and subsequently a high-quality aggregate representation. Interestingly, while random uniform patch sampling negatively impacts classification accuracy due to lower semantic content, it results in the highest positional equivariance as the model samples patches and corresponding saccade actions from a more diverse set of positions across the entire image, not just the salient regions. UMAP projections for PLS with and without action conditioning in Appendix B.5 further underscore the role of action conditioning in enabling positional equivariance.
Path Integration. In the context of eye movement-driven observations and in general sequential observations, another ability that naturally arises is path integration (McNaughton et al., 2006)—predicting the cumulative transformation/action from a sequence of actions. We evaluate this task in both eye movements in PLS (visual path integration) and object rotations in 3DIEBench (angular path integration). As shown in Figure 6, seq-JEPA demonstrates strong performance in both settings, with performance degrading gracefully as sequence length increases. Ablating action conditioning causes path integration to fail, whereas ablating the visual stream has only a minor impact—highlighting that action information is the dominant signal for this task. Full details of the path integration setup are provided in Appendix B.6.
@@
4.5 Ablations: Action Conditioning Is Key
To better understand the mechanisms underlying seq-JEPA’s invariant-equivariant representation learning and role of action conditioning, we perform a set of ablation experiments on 3DIEBench. Specifically, we study: (i) the role of action conditioning in the transformer and predictor; (ii) the impact of learning action embeddings and their dimensionality. Table 7 summarizes our ablation results. Removing action conditioning entirely causes a significant drop in equivariance ($R^{2}$ from 0.71 to 0.29), although classification accuracy remains high thanks to sequence aggregation and segregated invariance-equivariance in our model. Conditioning only the transformer or only the predictor leads to intermediate results, with predictor conditioning proving more critical for equivariance. We also vary the dimensionality of the learnable action embeddings: performance saturates around the default size of 128, with smaller sizes (e.g., 16 or 64) already sufficient to capture the rotation structure. Non-learnable embeddings significantly reduce equivariance, underscoring the importance of learned representations in the action pathway.

### v3

@@ Abstract
Joint-embedding self-supervised learning (SSL) commonly relies on transformations such as data augmentation and masking to learn visual representations, a task achieved by enforcing invariance or equivariance with respect to these transformations applied to two views of an image. This dominant two-view paradigm in SSL often limits the flexibility of learned representations for downstream adaptation by creating performance trade-offs between high-level invariance-demanding tasks such as image classification and more fine-grained equivariance-related tasks. In this work, we propose seq-JEPA, a world modeling framework that introduces architectural inductive biases into joint-embedding predictive architectures to resolve this trade-off. Without relying on dual equivariance predictors or loss terms, seq-JEPA simultaneously learns two architecturally separate representations for equivariance- and invariance-demanding tasks. To do so, our model processes short sequences of different views (observations) of inputs. Each encoded view is concatenated with an embedding of the relative transformation (action) that produces the next observation in the sequence. These view-action pairs are passed through a transformer encoder that outputs an aggregate representation. A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation. Empirically, seq-JEPA demonstrates strong performance on both equivariance- and invariance-demanding downstream tasks without sacrificing one for the other. Furthermore, it excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
@@ 1 Introduction
Equivariance is a crucial representational property for downstream tasks that require fine-grained distinctions. For example, given representations that are invariant to color, it is not possible to distinguish between certain species of flowers or birds (Lee et al., 2021; Xiao et al., 2021). Moreover, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024). While some equivariant SSL approaches have reported minor gains on tasks typically associated with invariance (e.g., classification) (Devillers and Lefort, 2022; Park et al., 2022; Gupta et al., 2023), a growing body of work highlights a fundamental trade-off between learning invariance and equivariance, i.e., models that capture equivariance-related style latents do not fare well in classification and vice versa (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024; Rusak et al., 2025). This trade-off has recently received theoretical support (Wang et al., 2024b), underscoring the need for new architectural and objective designs that can reconcile these competing goals.
@@
Our results demonstrate that individual encoded views in seq-JEPA become transformation/action-equivariant. Through ablations, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network. In contrast, the aggregate representation of views, produced at the output of the transformer proves highly effective for invariance-demanding downstream tasks. This emergent architectural disentanglement of invariance and equivariance is central to seq-JEPA’s competitive performance compared to invariant and equivariant SSL methods on both categories of tasks (Figure 3). Unlike most prior equivariant SSL methods (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), our model does not rely on explicitly crafted loss terms or objectives to achieve equivariance, nor is it instructed to learn the decomposition of two representations. Instead, the dual representation structure arises naturally from the model architecture and action-conditioned predictive learning.
Beyond resolving the invariance-equivariance trade-off, seq-JEPA further benefits from processing a sequence of observation views; we show that it performs well on tasks requiring integration over sequences of observations. In one scenario, inspired by embodied vision in primates, our model learns image representations without augmentations or masking, solely by predicting across simulated eye movements (saccades). In another setting, it performs path integration over sequences of actions—such as eye movements or 3D object rotations in 3DIEBench (Garrido et al., 2023). Our key contributions are as follows:
  • We introduce seq-JEPA, a self-supervised world model that learns architecturally distinct representations for invariance- and equivariance-demanding downstream tasks through sequential prediction over action-observation pairs, without requiring explicit equivariance losses or dual predictors.
@@ 2.2 Architecture
A backbone encoder, $f$ encodes the first $M$ views, producing representations $\{z_{i}\}_{i=1}^{M}$. Except for $z_{M}$, each $z_{i}$ is concatenated with its corresponding action embedding and passed to a transformer encoder $g$ (no MLP projector is used after the encoder), which aggregates the sequence of action-observation pairs. The transformer uses a learnable [AGG] token (analogous to the [CLS] token in ViT (Dosovitskiy et al., 2020)) to generate the aggregate representation:
@@
The ground truth $z_{M+1}$ is computed using a target encoder—an exponential moving average (EMA) of $f$. This target representation is passed through a stop-gradient operator (sg) to avoid representational collapse (Grill et al., 2020; Chen and He, 2021; Assran et al., 2023). The training objective is to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$ with the loss function:
@@
Figure 1: seq-JEPA is a self-supervised world model that leverages a sequence of action-observation pairs to learn architecturally distinct representations for downstream tasks requiring transformation invariance or equivariance.
@@ 2.3 Action and observation sets
To evaluate generalization across transformation types, we consider three sets of action-observation pairs (Figure 2). See Appendix A.1 for details of each setup.
3D Invariant Equivariant Benchmark (3DIEBench). The 3DIEBench dataset (Garrido et al., 2023) is designed to evaluate representational invariance and equivariance. It includes 3D object renderings with variations in rotation, floor hue, and lighting. In this benchmark, the action between two views is the relative difference of these three factors of variation. We primarily study equivariance to $SO(3)$ rotations and secondarily to appearance factors of floor and light hue.
Hand-Crafted Augmentations. In this setting, we use transformed views generated via common SSL augmentations (e.g., crop, color jitter, blur), and actions correspond to relative augmentation parameters. We use CIFAR100 and Tiny ImageNet for experiments in this setup, and follow EquiMod’s augmentation protocol (Devillers and Lefort, 2022).
Predictive Learning Across Saccades (PLS). Going beyond conventional transformations such as augmentations or 3D rotations, we show that seq-JEPA can learn visual representations from a sequence of partial observations thanks to architectural inductive biases—without relying on any hand-crafted augmentations. Our PLS has a similar flavor to I-JEPA (Assran et al., 2023) but does not require engineered masking strategies. In PLS, we train seq-JEPA on sequences of patches extracted from full-resolution images. For instance, with STL-10 dataset, we use small $32\times 32$ patches to form the observation sequence. In this setting, actions correspond to the relative positions between patch centers, simulating saccadic eye movements and inducing 2-D positional equivariance to representations. To select fixation points, we adopt two biologically inspired techniques that increase informativeness, reduce redundancy, and improve the downstream utility of the aggregate representation (Figure 2):
@@ 3.1 Compared methods and baselines
We compare seq-JEPA against both invariant and equivariant SSL baselines. Invariant methods include SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022) and VICReg with trajectory regularization (Wang et al., 2024a). Equivariant methods include SEN (Park et al., 2022), EquiMod (Devillers and Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024). For all baselines, architectural details are given in Appendix A.3. We also evaluate two hybrid baselines based on our architecture:
@@ 3.2 Training protocol
All models use ResNet-18 (He et al., 2016) as the backbone encoder. For action conditioning, we use a learnable linear projection to learn action embeddings (default action embedding is 128-d). The sequence aggregator in seq-JEPA is a lightweight transformer encoder (Vaswani et al., 2017) with three layers and four attention heads. The predictor is a 2-layer MLP with 1024 hidden units and ReLU activation. In order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we trained baselines that typically use MLP projectors in two variants: (1) with original MLP projector; and (2) with the MLP replaced by a transformer encoder and a sequence length of one. We did not see any benefit from switching to transformer projectors in any of the baselines, and include the transformer-projector results in Appendix B.2. All models are trained from scratch with a batch size of 512. We use 1000 epochs for 3DIEBench and 2000 epochs for other datasets to obtain asymptotic performance. We use AdamW for models with transformer projectors (including seq-JEPA) due to its stability and improved regularization in transformer training. For ConvNet-only models with MLP heads, we use the Adam optimizer. Full hyperparameters are detailed in Appendix A.
@@ 3.3 Evaluation metrics and protocol
To assess equivariance, we follow the protocol of Garrido et al. (2023) and train a regressor on frozen encoder representations to predict the relative transformation (action) between two views. We report the $R^{2}$ score on the test set. In addition to action decoding $R^{2}$, we also report retrieval-based metrics including Mean Reciprocal Rank (MRR), Hit@1, and Hit@5 (Kipf et al., 2020; Park et al., 2022; Garrido et al., 2023) to evaluate the quality of the predictor. As a proxy measure of invariance, we use top-1 classification accuracy of a linear probe on top of frozen representations. For all baselines, probes are trained on encoder outputs. For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 1) and report the number of observation views used during training and inference. For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 1) for seq-JEPA models in Appendix B.7. Training details of evaluation heads are given in Appendix A.2.
@@ 4.1 Quantitative evaluation on 3DIEBench
We use the 3DIEBench benchmark to quantitatively compare performance on equivariance- and invariance-demanding downstream tasks in seq-JEPA with baseline methods. This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object classification. Table 3 provides a summary of our evaluation on the 3DIEBench where equivariant methods have been conditioned on rotation (a 4-D quaternion representing the relative rotation between two views). In addition to the relative rotation between two views, in the last column we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view. For seq-JEPA, we trained models with varying training sequence lengths (denoted by $M_{tr}$ in the table) and measure the linear classification performance on top of aggregate representations with different inference lengths (denoted by $M_{val}$).
Among invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance. Adding a linear trajectory regularization loss to VICReg Wang et al. (2024a) without using ground-truth transformations improves over VICReg, which shows that imposing geometric priors can improve both invariant and equivariant performance—even when these priors do not fully materialize in the environment, e.g., when we have non-smooth angle changes as in 3DIEBench. Among equivariant baselines, SIE and ContextSSL yield strong rotation prediction performance due to their specialized equivariance predictors and loss functions, but underperform in classification. EquiMod and SEN offer better classification performance, yet compromise equivariance. In contrast, seq-JEPA achieves strong performance in both, matching the best rotation $R^{2}$ scores while exceeding baselines in classification, with gains increasing with inference sequence length $M_{val}$. Ablating action conditioning leads to a sharp drop in equivariance but retains classification accuracy, confirming our hypothesis that action-conditioned sequential aggregation enables a distinct representational structure for invariance- and equivariance-demanind tasks. For additional results on 3DIEBench including MRR, Hit@1, and Hit@5 metrics, performance of models with varying training and inference sequence lengths and models conditioned on both rotation and color, and evaluation on an out-of-distribution (OOD) set of 3DIEBench see Appendices B.1 to B.6.
@@
Table 1: Evaluation on 3DIEBench for linear probe classification (invariance-demanding) and rotation prediction (equivariance-demanding). Equivariant models and seq-JEPA are conditioned on rotation. For seq-JEPA, training and inference sequence lengths are denoted by $M_{tr}$ and $M_{val}$.
@@ 4.3 Evaluation with Hand-Crafted Augmentations
We assess downstream performance under hand-crafted augmentations by training on CIFAR100 and Tiny ImageNet (Table 2). Models with action conditioning are trained conditioned on crop, color jitter, blur, or all three (indicated in the first column of the table). seq-JEPA consistently achieves higher equivariance than both invariant and equivariant baselines across all transformations. Notably, except for the model trained on CIFAR-100 and conditioned on blur, the best equivariance performance for a given augmentation is achieved when the model is specialized and conditioned only on that augmentation. Furthermore, ablating actions (last row in the table) causes seq-JEPA to lose its equivariance across transformations compared to action-conditioned models. Overall, our model outperforms both invariant and equivariant families in terms of equivariance, while being competitive in terms of classification performance. For additional results with varying training and inference sequence lengths, see Appendix B.6.
@@ 4.4 Predictive Learning across Saccades and Path Integration
In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations to learn visual representations. In Table 6, seq-JEPA reaches 83.44% top-1 accuracy on STL-10, comparable to SimCLR (85.23%) trained with full-resolution images and strong augmentations. This gap narrows further when increasing the inference sequence length from $M_{val}=4$ to $M_{val}=6$. Ablating action conditioning causes the accuracy on top of the aggregate representations to drop sharply, indicating that 2-D positional awareness is essential to forming semantic representations across simulated eye movements. Compared to Conv-JEPA—which accumulates prediction losses pairwise—seq-JEPA performs better in classification, highlighting the importance of sequence aggregation when dealing with partial observations in SSL. Further ablations show that saliency-driven sampling and IoR are critical for forming informative, non-overlapping patch sequences and subsequently a high-quality aggregate representation. Interestingly, while random uniform patch sampling negatively impacts classification accuracy due to lower semantic content, it results in the highest positional equivariance as the model samples patches and corresponding saccade actions from a more diverse set of positions across the entire image, not just the salient regions. The UMAP projections for PLS with and without action conditioning in Appendix B.8 further underscore the role of action conditioning in enabling positional equivariance.
Path Integration. In the context of eye movement-driven or any sequential observations, an ability that naturally arises from predictive learning is path integration (McNaughton et al., 2006), i.e., predicting the cumulative transformation/action from a sequence of actions. We evaluate this task in both eye movements in PLS (visual path integration) and object rotations in 3DIEBench (angular path integration). As shown in Figure 6, seq-JEPA demonstrates strong performance in both settings, with performance degrading gracefully as sequence length increases. Ablating action conditioning causes path integration to fail, whereas ablating the visual stream has only a minor impact—highlighting that action information is the dominant signal for this task. Full details of the path integration setup are provided in Appendix B.9.
@@
4.5 Action Conditioning Ablations
To better understand the mechanisms underlying seq-JEPA’s invariant-equivariant representation learning and role of action conditioning, we perform a set of ablation experiments on 3DIEBench. Specifically, we study: (i) the role of action conditioning in the transformer and predictor; (ii) the impact of action embedding dimensionality. Table 7 summarizes our ablation results. Removing action conditioning entirely causes a significant drop in equivariance ($R^{2}$ from 0.71 to 0.29), although classification accuracy remains high thanks to sequence aggregation and segregated invariance-equivariance in our model. Conditioning only the transformer or only the predictor leads to intermediate results, with predictor conditioning proving more critical for equivariance. We also vary the dimensionality of the learnable action embeddings: performance saturates around the default size of 128, with smaller sizes (e.g., 16 or 64) already sufficient to capture the rotation structure.